# Large Language Models:

### v1

@@ Abstract
Large language models (LLMs) are being used in economics research to form predictions, label text, simulate human responses, generate hypotheses, and even produce data for times and places where such data don’t exist. While these uses are creative, are they valid? When can we abstract away from the inner workings of an LLM and simply rely on their outputs? We develop an econometric framework to answer this question. Our framework distinguishes between two types of empirical tasks. Using LLM outputs for prediction problems (including hypothesis generation) is valid under one condition: no “leakage” between the LLM’s training dataset and the researcher’s sample. Using LLM outputs for estimation problems to automate the measurement of some economic concept (expressed by some text or from human subjects) requires an additional assumption: LLM outputs must be as good as the gold standard measurements they replace. Otherwise estimates can be biased, even if LLM outputs are highly accurate but not perfectly so. We document the extent to which these conditions are violated and the implications for research findings in illustrative applications to finance and political economy. We also provide guidance to empirical researchers. The only way to ensure no training leakage is to use open-source LLMs with documented training data and published weights. The only way to deal with LLM measurement error is to collect validation data and model the error structure. A corollary is that if such conditions can’t be met for a candidate LLM application, our strong advice is: don’t.
@@ 1 Introduction
How should large language models (LLMs) be used in economics research? [^1] Good reviews of machine learning more broadly and its applications within economics include Varian (2014); Mullainathan and Spiess (2017); Athey (2018); Gentzkow, Kelly and Taddy (2019); Dell (2024). Since they are easy-to-use, general-purpose tools, researchers are already using LLM outputs to form predictions, label text, answer surveys, simulate humans in experiments, generate hypotheses, and even produce data for time periods and places for which such data don’t currently exist. [^2] For example, generating data on, say, inflation expectations or consumer sentiment during pre-Revolutionary America. There are many adjacent uses which we do not consider here, such as the use of LLMs as computational tools like others that occupy our workflow: search engines, and coding copilots. Korinek (2023, 2024) provide a useful description of all the ways LLMs can be used in this way. Our focus instead is on the uses of LLMs not in day-to-day workflow but directly in empirical analysis. These uses are no doubt creative, but are they valid? How can we assess whether empirical research using LLM outputs is leading to correct inferences?
Consider how we normally answer this question for a more familiar econometric procedure like ordinary least squares (OLS). At its core, OLS is nothing more than an algorithm that, when applied to a dataset, returns the coefficients $\widehat{\beta}$ that minimize the sum of squared residuals. Econometrics clarifies what assumptions the researcher must believe about the underlying data generating process (DGP) to justify possible interpretations of the algorithm’s output. If we want to interpret $\widehat{\beta}$ as the best linear unbiased estimator of the conditional expectation function, the Gauss-Markov theorem tells us what we need to be willing to assume (and defend) about the DGP; if we want to interpret $\widehat{\beta}$ causally, we need to be willing to defend the conditional independence assumption. In this sense, econometrics provides empirical researchers with the terms of a contract: to endow an estimate with a particular interpretation, these are the assumptions you must defend.
Appropriately using LLM outputs requires the same sort of contract: what assumptions about LLMs must be defended in order to draw some particular inference? Yet at present, no such contracts exist. In fact, producing such a contract is challenging given how econometrics typically models data procedures. Our contracts for OLS are based on knowing the exact operations of OLS on a data matrix. In contrast, LLMs are extraordinarily complex machine learning models involving many layers of interactivity and billions of parameters. Their training datasets and architecture (among many other details) are proprietary—that is, intentionally hidden from the user. [^3] As an example, the GPT-4o technical report writes “this report contains no further details about the architecture (including model size), hardware, training compute, dataset construction, training method, or similar” (OpenAI, 2023). They are also a diverse, dynamic set of commercial products. Different families of LLMs make different choices along each of these dimensions. The state-of-the-art is ever-evolving: new LLMs are released constantly. Deriving guarantees from the exact operations of these models, like we do for OLS, is simply not feasible.
The goal of this paper is to provide an econometric framework that provides valid LLM contracts despite the complexity of these models. Our framework considers settings in which a researcher uses an LLM to process text and produce outputs that can be related to traditional economic variables in downstream analyses. Precisely because these algorithms are opaque and varied, we treat everything about how the LLM goes from its training data to a particular text response as a black box. The framework distinguishes between two types of empirical applications — prediction and estimation — and provides four main results.
Our first result relates to prediction problems, in which the researcher predicts a linked economic variable using the associated text. For example, in asset pricing, we would like to know how well stock prices can be predicted using the text of news headlines. We show that for LLM outputs to be validly used in prediction problems, the LLM must satisfy a single condition we refer to as “no training leakage.” Suppose the researcher prompts the LLM on each collected piece of text to form predictions, and evaluates the quality of the LLM’s resulting predictions by calculating its sample average loss on the researcher’s dataset. We show that the LLM’s sample average loss only reflects its true out-of-sample predictive performance if and only if there is no training leakage, meaning there is no overlap between the text in the LLM’s training dataset and the researcher’s own dataset. Intuitively, in this workflow, the researcher is using the LLM as-if it were some prediction function from text to the economic variable, and is then using their own collected dataset as-if it were a test sample. This is only valid if the LLM has not been trained on the test sample.
No training leakage is often violated in naive uses of LLMs precisely because their training datasets are both immense and intentionally obscured to researchers. This is a well-known challenge in computer science, where researchers worry that LLMs have increasingly been trained on the benchmark datasets commonly used to evaluate LLM performance. We present new empirical results showing that the no-training-leakage condition is also violated in settings relevant to economists, through two applications to finance and political economy. We show that GPT-4o has likely been trained on economic datasets, since it appears to have exactly memorized the text of many financial headlines and Congressional bills. [^4] In recent work, Sarkar and Vafa (2024) show LLMs used to predict a future outcome from some text can inadvertently draw on information from the time of the outcome rather than the time of the text. The authors refer to this as “look-ahead bias,” which violates our “no training leakage” condition.
Even though it can be violated, we argue that “no training leakage” can serve as a credible assumption in many applications since it can be controlled by the researcher’s choice of large language model. But that requires using open-source LLMs that clearly document their underlying training data and/or provide a clear time-stamp beyond which the LLM has not been updated, such as the Llama family of models (Touvron et al., 2023; Dubey et al., 2024) or the StoriesLM family (Sarkar, 2024). Researchers can thereby mechanically enforce that the no-training-leakage condition is satisfied since it is known that these models have not been trained on any strings past a documented date.
Our second result relates to estimation problems, in which the researcher wants to measure some economic concept (e.g., positive or negative financial news, hawkish attitudes, or policy topic) expressed by some text in order to estimate some downstream economic parameter. For example, in studying partisanship, how does the policy topic of a Congressional bill relate to the ideology of its sponsor? The researcher has access to a resource-intensive, “gold-standard” process for measuring the economic concept. For instance, with enough time and effort, the text could be reliably labeled for the concept by the researcher or trained staff. Because the labeling of the text is often prohibitively costly or time-consuming, the researcher hopes to automate this process—to use the LLM to label text for the concepts.
In order for an LLM to successfully automate the existing measurement process, we demonstrate the key condition is that there must be no measurement error in the LLM’s outputs. The plug-in estimate using the LLM’s labels recovers the target estimate if and only if the LLM’s labels reproduce the existing gold-standard measurement process.
This no-measurement-error assumption is simply not credible. A rapidly growing body of research documents the brittleness of LLMs across computer science benchmarks — what seem like small modifications to a task (choice of LLM or specific prompt, for example) lead to substantial changes in LLM performance. We present new results showing the implications of this brittleness for downstream estimation of economic parameters. For the applications of relating financial news headlines to stock prices, and relating the policy topic of a Congressional bill to characteristics of the bill’s sponsor, we show that variability in the LLM’s outputs across models and prompts leads to large changes in magnitude, statistical significance and even sign of resulting regression coefficients.
Our third result is a constructive solution for how to use LLMs for estimation problems in light of these challenges. Since LLMs are brittle and therefore not a general-purpose technology for producing economic labels, researchers will need some context-dependent playbook for using them in estimation problems. Simply calculating overall accuracy measures of the LLM labels is not sufficient because it is not just the magnitude of the measurement error that matters (sometimes reported in existing studies), but also its covariance with the other economic variables in the regression (unfortunately rarely if ever reported). Precisely because we have little insight into their design, it is difficult to articulate reasonable statistical assumptions on the properties of the LLM’s measurement error.
We therefore borrow from an idea long known in economics: the only solution to measurement error is to collect validation data containing gold standard measurements and explicitly model the LLM’s measurement error. While this is well-known in labor economics (e.g., Bound and Krueger, 1991; Bound et al., 1994; Bound, Brown and Mathiowetz, 2001) and well-studied in econometrics (e.g., Chen, Hong and Tamer, 2005; Schennach, 2016), it has only recently been revived in machine learning, such as Wang, McCormick and Leek (2020); Angelopoulos et al. (2023); Egami et al. (2024). We illustrate the value of validation data in the context of linear regression, using the LLM’s output as either a covariate or dependent variable. Through asymptotic arguments, we show that debiasing LLM outputs using a validation sample preserves our usual econometric guarantees of consistency and asymptotic normality, while still drastically reducing the costs of measuring the economic concept of interest. We empirically demonstrate this performance in finite samples with Monte Carlo simulations using data on Congressional bills.
The implication is that for estimation problems, LLM outputs are not substitutes for gold standard labels. When done correctly, LLM outputs instead serve to amplify a small validation sample, enabling the researcher to draw correct inferences at a lower cost. A corollary is that researchers should not use LLM outputs for estimation problems in applications where such validation data cannot be collected.
Our final result is to show that this framework is powerful enough to help economists carry out valid inference not just for familiar empirical tasks like prediction and estimation, but for more novel ones as well. For example, we argue that the use of LLMs for hypothesis generation has the same structure as other prediction problems. That means our framework can clarify what assumptions are required for the valid use of LLMs for this purpose—no training leakage. Or consider the use of LLMs to simulate human subject responses. We note the similarity of this application to other estimation problems, and hence the vital importance of collecting validation data (e.g., running experiments, collecting surveys, etc. from at least some real subjects).
Notice that none of these results required articulating how exactly the LLM goes from its training data to its output. Black-boxing new AI tools in this manner might initially seem unsatisfying, if not unnerving. After all, isn’t it important to understand the trick of, say, positional encoding and all the innards of the transformer architecture? Economists, however, already black-box important upstream elements of familiar econometric procedures — we just often don’t recognize it. For example, few researchers could describe the specific numerical implementation of OLS in their preferred statistical software (in fact many of us have not realized these implementations differ across software programs). We are nonetheless happy to proceed with the returned estimate $\widehat{\beta}$ so long as we know they satisfy a key property, in this case of having minimized the sum of squared residuals. Our empirical framework provides a similar set of properties for LLMs that allows us to use their outputs in empirical research.
LLMs represent a remarkable new technological advance. They purport to be a general-purpose technology, one that is both a substitute for intelligence and a source of ground truth. Their impressive performance on certain activities—as well as a great deal of hype—lures us into treating their output as if it were indeed ground truth. Their general-purpose makes us feel we should be able to apply them everywhere. But those temptations, understandable as they may be, must be resisted.
Our over-arching argument is that there shouldn’t be a separate “VIP lane” for LLMs into our economics research. The bridge from statistical output to conclusions about the world should be carefully guarded. The cardinal sin of empirical research is to use a data procedure without clearly stating our assumptions and defending them. No one would claim that some instrumental variables estimate is a local average treatment effect without defending the monotonicity assumption (Imbens and Angrist, 1994; Angrist, Imbens and Rubin, 1996), or present some panel-data regression without defending its standard error calculation (Bertrand, Duflo and Mullainathan, 2004) - much less fail to present a standard error at all. Yet analogous practices are rampant with LLMs. While LLMs seem different, just like any other procedure in our empirical toolkit, they require rigorous econometric contracts clarifying when and how economists can make valid inferences based on them.
2 Evaluating General-Purpose Technologies: the LLM Conundrum
Why has our research community been willing to create a VIP lane for LLMs straight into our studies? For two reasons. The first relates to the impressive performance of LLMs on benchmark evaluations, combined with memorable examples of their capabilities. The second is our tendency to draw inferences and generalize from how LLMs performed on some tasks to how they are likely to perform on other tasks. After all, if we saw a human ace the math GRE, we would naturally assume this person would perform well on all manner of other math problems as well. We tend to make the same assumption about LLMs.
Yet both reasons turn out to be problematic. The domains on which LLMs perform well turn out to be quite selective, especially when compared to the grand ambition to have LLMs serve as general-purpose technologies. And the way that the capabilities of LLMs generalize across tasks turns out to be quite different than what we intuitively assume. In what follows we give a sense for how truly odd the behavior of LLMs can be, and consequently how far they are from being able to perfectly substitute for human intelligence, much less serve as a sort of “super intelligence.”
2.1 Benchmark Evaluations and Anthropomorphic Generalization
The central difference between LLMs and supervised learning algorithms is that LLMs aspire to serve as useful tools not just for a single task, but all tasks — to be a general-purpose technology (GPT). The scope of that ambition is both the most remarkable feature of LLMs and also their Achilles’ heel.
Because supervised learning algorithms (e.g., convolutional networks) aim to tackle a single task (e.g., image classification), it is straightforward to evaluate any algorithm’s prediction — to know how well an algorithm works and which alternative design works better. This was operationalized by the machine learning community through the creation of open public prediction tasks that competing algorithms could be deployed on to assess performance, the so-called common task framework (Donoho, 2024). The resulting competitions (e.g., the Netflix prize, ImageNet, etc.) helped spur innovation and progress.
But there is an obvious challenge in extending the common task framework to any new tool (like LLMs) that aspires to serve as a GPT: how does one rank-order the performance of competing LLMs when they aim to be useful on any task? How do we develop open common-task competitions to evaluate LLMs when there is no common task?
The current solution is to simply try building bigger and more diverse benchmark evaluations. For example, the “Beyond-the-Imitation-Game benchmark” (BIG-bench) collects problems on 204 tasks ranging from math and analogical reasoning questions to reading comprehension and social reasoning problems (Srivastava et al., 2022). The “Massive Multitask Language Understanding” (MMLU) benchmarks collect questions across 57 different scientific and humanistic disciplines (Hendrycks et al., 2020). Even more common is the evaluation of LLMs by their designers on standardized exams that aim to test general-purpose knowledge in people, such as the SAT, the GRE, and AP exams. If the collection of questions across these benchmark evaluations were actually what users ultimately wanted to deploy LLMs on, then comparing performance across language models could be reduced to constructing task performance metrics on these benchmarks.
However, we do not intrinsically care about these benchmark tasks — after all, relatively few researchers will ever deploy an LLM to take the SAT or answer abstract reasoning puzzles. Rather, we use these benchmark evaluations to generalize about the capabilities of LLMs on new tasks. This reflects a deeper bias we have when we engage with LLMs: we tend to engage in anthropomorphic generalization. Since it is difficult to imagine a person who could accomplish these feats yet fail on related (or even simpler) tasks, we tend to generalize the performance of an LLM onto new tasks the same way we would generalize human performance across tasks. Surely a large language model that can write a proof that there are infinitely many primes in the form of a poem (Bubeck et al., 2023) must be able to label text or answer surveys in economics research.
Evidence to this effect comes from vafa2024humangeneralization, who present online subjects with pairs of questions and ask subjects to predict whether a human (or algorithm) would get the second question correct depending on how they did with the first question. The authors refer to a person’s ability to accomplish this task as the “human generalization function.” While the human generalization function predicts the performance of other humans well, it is misaligned with the performance of LLMs — these models often get questions wrong that we expect them to get right based on our observations of their performance on previous tasks. DreyfussRaux(24) document a similar phenomenon, showing that users extrapolate the performance of LLMs based on an intuitive notion of task difficulty and explore implications for LLM usage. [^5] These findings are also backed up by experimental evidence from giving large language models to managers and consultants to solve different tasks, alternating between tasks on which the models do well and other tasks of seemingly-similar difficulty levels that the models handle poorly—what the authors call AI’s “jagged technological frontier” (Dellacqua(23)).
This research, in other words, confirms that people engage in a naive heuristic of anthropomorphic generalization, which is surely one important reason we are all so willing to treat the labels of LLMs as ground-truth and plug them into our research pipelines.
2.2 Impressive Feats, But Also Puzzling Examples
This anthropomorphic generalization of the LLM’s capacity is common, understandable (given their seemingly human interface), and deeply problematic. Recent work in computer science shows the capabilities of LLMs are brittle: For every example of impressive performance, there is a counterexample that leaves users scratching their heads.
The type of brittleness that has attracted the most public attention is the periodic tendency of LLMs to hallucinate—to report back plausible-sounding “facts” that are not actually facts. For example, an LLM asked to provide proof that dinosaurs created a civilization will reply that there is fossil evidence of dinosaur tools, and even credit dinosaurs with the invention of primitive art forms like stone engravings (Szempruch(2023)). This tendency to hallucinate is not limited to trivial application domains, as was discovered by the lawyer who used an LLM to prepare a legal brief, only to discover it cited a number of “cases” that are not real cases (Bohannon(2023)). These hallucinations are not rare events; one study found that LLMs hallucinate in $58\%$ of legal applications (Dahl_2024). While some may hope that there exists some future technological solution to hallucination, recent work provides statistical and computational arguments suggesting this may be inevitable — an intrinsic and unavailable feature of the language generation problem (e.g., kalai2024calibratedlanguagemodelshallucinate; xu2024hallucinationinevitableinnatelimitation).
While hallucinations are well-known, economists and other users who are impressed by some LLM’s performance on the math SAT or math Olympiad might be surprised to see examples of how poorly these same LLMs can perform on other math problems. For example, the same LLM that can reliably solve (9/5) x + 32 cannot solve (7/5) x + 31; an LLM that can reliably implement common ciphers such as shift by 13 cannot implement other variations (mccoy2023embersautoregressionunderstandinglarge).
The performance of LLMs turns out to be remarkably sensitive to seemingly minor details. An LLM that can correctly answer a given multiple choice question will often answer the same question incorrectly after the order of the answers has been permuted (zong2024foolvisionandlanguage). LLMs struggle on “counterfactual” versions of tasks — for example, being able to program a list sort in 0-based indexing but unable to perform the same task in 1-based indexing (wu2024reasoningrecitingexploringcapabilities). Similar sensitivity has been documented on logical reasoning tasks (lewis2024usingcounterfactualtasksevaluate). Their behavior can be substantially influenced by appending an adversarial string to an existing question (zou2023universaltransferableadversarialattacks). The same GPT-4 model that demonstrates impressive spatial awareness in describing how to carefully stack a book, a laptop, nine eggs, a bottle and a nail does very poorly for stacking a pudding, a marshmallow, a toothpick and a glass of water (MelanieMitchell(23)-Brittleness).
Despite their many amazing feats and human-like interface, LLMs do not reason the way humans reason. An LLM trained on “A is B” will not know “B is A.” For example, an LLM trained on the fact that “Tom Cruise’s mother is Mary Lee Pfeiffer” cannot answer “Who is Mary Lee Pfeiffer’s son?” (berglund2023reversal). LLMs struggle on variations of the question: “Alice has $N$ brothers and she also has $M$ sisters. How many sisters does Alice’s brother have?” (nezhurina2024alicewonderlandsimpletasks). When asked how to ferry a single farmer and a single sheep from one side of a river to the other using a boat with enough room for one person and one animal, the LLM claims this requires at least three trips (e.g., Fraser(2024)-Claude; Fraser(2024)-GPT).
Despite this accumulating body of evidence, outputs from these tools are being uncritically incorporated into a growing number of empirical studies. This sets up the problem statement that we seek to address. Since LLMs are easy-to-use yet brittle tools for processing text, under what conditions can researchers safely plug-in the outputs of LLMs into a research pipeline? Our argument, in short, is that “I played around with it and it seemed to do well” is not a valid justification for widespread use of such models within economics research. To provide a more constructive alternative, we develop next an econometric framework to explore this question formally.
3 An Empirical Framework for LLMs
A key feature of modern LLMs is that we typically interact with them as black boxes: even though key choices in their design and the exact contents of their training datasets are unknown to us, we nonetheless use them to generate responses given prompts. We are willing to use these LLMs as a generic tool for processing almost any text—a true general-purpose technology—given their impressive performance on benchmark evaluations. But does this heuristic make any sense?
In this section, we develop a framework that helps clarify what properties an LLM must satisfy in order for us to safely incorporate it into empirical research. While social scientists in practice tend to treat LLMs as an all-purpose black box, in reality the model’s performance will hinge critically—as our framework demonstrates—on what text the algorithm is trained on, to what text we seek to apply it, and for what purpose.

### v3

@@ Abstract
Large language models (LLMs) enable researchers to analyze text at unprecedented scale and minimal cost. Researchers can now revisit old questions and tackle novel ones with rich data. We provide an econometric framework for realizing this potential in two empirical uses. For prediction problems – forecasting outcomes from text – valid conclusions require “no training leakage” between the LLM’s training data and the researcher’s sample, which can be enforced through careful model choice and research design. For estimation problems – automating the measurement of economic concepts for downstream analysis – valid downstream inference requires combining LLM outputs with a small validation sample to deliver consistent and precise estimates. Absent a validation sample, researchers cannot assess possible errors in LLM outputs, and consequently seemingly innocuous choices (which model, which prompt) can produce dramatically different parameter estimates. When used appropriately, LLMs are powerful tools that can expand the frontier of empirical economics.
JEL Codes: C10, C52.
Keywords: Large language models, prediction, estimation, leakage, validation.
@@ 1 Introduction
Large language models (LLMs) enable economists to process text at unprecedented scale and minimal cost, unlocking questions previously out of reach: predicting stock returns from earnings call transcripts, measuring partisan polarization in social media posts, backcasting historical consumer sentiment from newspapers, or simulating survey responses cheaply. By making these questions – and countless others – tractable, LLMs offer transformative potential for empirical research.
To realize this potential, we must answer a practical question: how should LLM outputs be incorporated into empirical workflows? Can they be plugged in “as-is,” or do our estimation strategies require adjustment to use these powerful tools?
This question is challenging because LLMs resist traditional statistical analysis. LLMs are complex, often proprietary, constantly evolving, and trained on sprawling, heterogeneous corpora that defy tractable modeling – extraordinary engineering achievements accomplished without our usual statistical foundations. Existing evaluation methods have been remarkably effective for developing better models but offer limited guidance for how any given LLM will perform on a new task.
We develop an econometric framework that addresses these complexities and provides practical guidance for empirical research using LLM outputs, focusing on two empirical uses: prediction problems and estimation problems. For each use, we clarify what assumptions and practices ensure valid conclusions.
In prediction problems, researchers use collected text to predict some economic outcome (e.g., predicting stock prices from corporate earnings calls). Extracting meaning from text requires modeling the complex structure of language. LLMs, having already learned this structure from enormous training corpora, can possibly serve as the foundation upon which economists can build in prediction problems — either by directly prompting an LLM to make a prediction or by using its representations as features in a predictor.
Using LLMs in prediction problems requires one condition: “no training leakage.” If a researcher forms predictions using an LLM and evaluates those predictions on an evaluation dataset, this reflects the LLM’s out-of-sample performance if and only if there is no overlap between the LLM’s training dataset and the researcher’s dataset. This can be violated because many LLMs are trained on intentionally obscured datasets. Computer scientists have found that LLMs are often trained on common benchmark evaluations. We find that LLMs appear to be trained on economic datasets; see also glasserman2023assessinglookaheadbiasstock; sarkar2024lookahead; lopezlira2025memorizationproblemtrustllms among others.
We provide guidance on enforcing no training leakage, clarifying that it requires careful attention to both the choice of model and research design. For example, when the goal is to predict future, unseen documents (such as tomorrow’s financial news), open-source LLMs with fixed, published weights (e.g., touvron2023llama2openfoundation; dubey2024llama3herdmodels) or time-stamped training data (e.g., Sarkar(24); he2025chronologicallyconsistentlargelanguage) can be paired with evaluation samples constructed only from documents published after the model’s publication date or training cutoff.
In estimation problems, researchers estimate relationships between economic concepts expressed in text (e.g., partisanship in social media posts) and downstream parameters (e.g., the causal effect of a policy change). There is some resource-intensive procedure for measuring the economic concept; if it could be scaled, we would use these measurements and report the resulting estimates. Often, this would involve the researcher carefully reading each text piece themselves. We would like to instead use an LLM to economize on measurement costs. How can we use LLM outputs for valid downstream inference?
We highlight a constructive solution borrowing from an old idea in econometrics: collect a small validation sample and use it to empirically correct for possible LLM errors. We illustrate how researchers can incorporate validation data in the context of linear regression. Debiasing LLM outputs preserves our usual econometric guarantees of consistency and asymptotic normality. This has been well-studied in econometrics (e.g., LeeSepanski(95)-ErrorsInVariable; chen2005measurement; schennach2016recent) and extended in machine learning, such as WangMcCormickLeek(20), AngelopoulosEtAl(23)-PPI, EgamiEtAl(24)-ImperfectSurrogates, BattagliaEtAl(24), and carlson2025unifyingframeworkrobustefficient. We illustrate that incorporating debiased LLM outputs can substantially improve the precision of downstream estimates – resulting in estimates that are often more precise than using validation data alone. Used correctly, imperfect LLM outputs serve not as substitutes but as amplifiers of validation samples, allowing researchers to achieve tighter standard errors.
Absent a validation sample, researchers cannot assess the magnitude or pattern of errors in LLM outputs—and therefore cannot evaluate their impact on downstream parameter estimates. We demonstrate this problem empirically: absent a validation sample, seemingly innocuous choices—which LLM to use, how to phrase the prompt—lead to dramatically different parameter estimates in applications to finance and political economy, with coefficients varying in magnitude, sign, and significance. In estimation problems, researchers have no choice but to invest effort and collect validation samples when using LLMs.
Finally, this framework is flexible enough to account for creative uses of LLMs in economics. We argue that using LLMs for hypothesis generation (e.g., ludwig2024machine) can be cast as a form of prediction problem, so that the key assumption required is again no training leakage. LLM simulation of human subjects in surveys or experiments (e.g., park2023generativeagentsinteractivesimulacra; horton2023large; manning2024automatedsocialsciencelanguage; park2024generativeagentsimulations1000) can be thought of as an estimation problem, and so in-silico subjects can amplify – but not fully replace – a small validation sample of actual human responses.
Our framework clarifies when and how researchers can harness LLMs in empirical research. The requirements are straightforward: ensure no training leakage for prediction and collect small validation samples for estimation. We provide a checklist based on our framework in Section 6. These simple practices unlock the transformative potential of LLMs for empirical research.
2 An Econometric Framework for Large Language Models
LLMs are complex architectures trained on vast corpora through multi-stage processes: pretraining (learning to predict next tokens), instruction fine-tuning, reinforcement learning from human feedback (RLHF), and reinforcement learning from verifiable rewards (RLVR). Reasoning models employ test-time computation to further improve performance. The field evolves rapidly, with new architectures, training procedures, and capabilities emerging regularly. [^1] Many resources are available about the technical foundations of large language models. See, for example, chang2024_llmevals, minaee2025largelanguagemodelssurvey, and zhao2025surveylargelanguagemodels. For overviews aimed at economists, see Korinek(23), dell2024deeplearningeconomists, Korinek(24), and ash2024llm.
This complexity creates a fundamental challenge for econometric analysis. We typically study statistical procedures by articulating assumptions about the data-generating process and combining them with a computational understanding of the procedure. This approach is at present intractable for LLMs. They have billions of parameters, proprietary training datasets, and algorithms that resist formal characterization.
We adopt a different strategy based on how economists actually use these tools. We treat LLMs as black boxes – prompting them or extracting embeddings without characterizing internal mechanisms – and identify conditions they must satisfy for valid empirical use. We focus on two applications: prediction problems and estimation problems. By black-boxing their inner workings, our approach provides guidance that is robust to inevitable changes in architectures and training procedures.